A sample of the rul'd pt that has an input of 7.5
Answer #1

Part-A : The half life of this nuclide will be given as -

using a formula, we have

N = N0 e-\lambdat

(9.45 x 1010 Bq) = (7.54 x 1011 Bq) e-\lambdat

[(9.45 x 1010 Bq) / (7.54 x 1011 Bq)] = e-\lambdat

(0.12533156) = e-\lambdat

-\lambda t = ln (0.12533156)

\lambda = (2.07679) / (92.4 min)

\lambda = 0.02247 min-1

We know that,     \lambda = (0.693) / t1/2

t1/2 = (0.693) / (0.02247 min-1)

t1/2 = 30.8 min